Waddle Company amended its defined benefit pension plan on January 1, 2024, to increase retirement benefits earned with each service year. The actuary estimated the prior service cost to be $216,000. Waddle's 80 present employees are expected to retire at the rate of about 10 each year at the end of each of the next eight years.

Required:

  1. Using the service method, calculate the amount of prior service cost to be amortized to pension expense in 2024.

  2. Using the straight-line method, calculate the amount of prior service cost to be amortized to pension expense in 2024.
Answer #1

2024-2031

80(year 2024)+70+60+50+40+30+20+10(yr 2023)= 360

1) 216,000*80/360=48,000 answer


2) 216,000/8yrs= 27,000 answer

answered by: anonymous